Find the multiplicative inverse of 2.25

ok..I get that multiplicative inverse of 7 is 1/7 and 2/3 is 3/2...but I do not for some reasom get the the multiplicative inverse of 2.25!!!! Please explain thanks
: Find the multiplicative inverse of 2.25

ok..I get that multiplicative inverse of 7 is 1/7 and 2/3 is 3/2...but I do not for some reasom get the the multiplicative inverse of 2.25!!!! Please explain thanks

Answer by Earlsdon(3748)

Find the multiplicative inverse of 2.25
You will first need to convert the decimal number 2.25 to an improper fraction.
2.25 = 2 1/4 = 9/4
The multiplicative inverse of 9/4 is 4/9, so...
The multiplicative inverse of 2.25 = 4/9
However, if you need your answer to be in decimal form, then...
4/9 = 0.444...
The multiplicative inverse of 2.25 = 0.444...